OpenStudy (anonymous):

What number is 175% of 76?

OpenStudy (shamil98):

1.75x = 0.76

OpenStudy (kainui):

What's 100% of 10? What's 50% of 10? You already know the answer is 10 for the first question and 5 for the second question I asked, right? So let's try to figure out how to get there. So when we have 100% we multiplied 10 by 1. When we got 5, we multiplied 10 by .5 to get 5. So now we sort of see a pattern. You can take your percentage and move the decimal over 2 places to the left and multiply that decimal by your number. If you ever forget, just think of simple examples like this to figure it out again! Easy peasy!

OpenStudy (shamil98):

or is it 0.76x =1.75 im prob wrong

OpenStudy (shamil98):

oh. 1.75x = 76

OpenStudy (shamil98):

my brain is turned off atm

OpenStudy (mathlegend):

\[\frac{ is }{ of } = \frac{ Percent }{ 100 }\]

OpenStudy (mathlegend):

\[\frac{ x }{ 76 } = \frac{ 176 }{ 100 }\]

OpenStudy (mathlegend):

Can you solve the rest @spanntasticfam ?

OpenStudy (anonymous):

Thank you so much! Yes I have it now! :)

hero (hero):

It also works to do the literal interpretation: \[\frac{175}{100} \times 76\] Reduce, simplify and multiply

hero (hero):

\[\frac{(25)(7)}{(25)(4)}(76)\]\[=\frac{7}{4}(76)\]\[=7(19)\]

OpenStudy (mathlegend):

Oh, pardon me, my friend... @spanntasticfam make sure you put 175%... I accidentally put 176%.